1. A negative point charge of 10-6 coulomb is situated in air at the origin of a rectangular coordinate system. A second negative point charge of 10-4 is situated on the positive x-axis at a distance of 50 cm from the origin. What is the force on the second charge due to first?
1
Expert's answer
2012-03-13T07:33:16-0400
Let's provide some denotations:
q1 = - 10^(-6) - the charge of the first point
q2 = - 10^(-4) - the charge of the second point

Let's use the Coulomb's law now:

F = k*q1*q2/r² = 8.9877*(-10^(-6))*(-10^(-4))/0.5² = 8.9877*10^(-10)/0.25 = 39.9508*10^(-10) N.